6. 10 Pts Find the general solution of the given higher-order differential equation y (4) - 2y - 8y = 0
Answer #1

That's easy.

Have a great day!!!

sol! y4_2y-ey=0 :(04_202-8)y=0 : The auxiliary eq” is mia2m²8=0 ame_4m²72²8=0 : m2 (m²-4)+2(424)=0 3442+2) (m? 4) = 0 ... m=